OpenStudy (anonymous):

Solve the given equation. tan θ − 3 cot θ = 0

OpenStudy (ghazi):

\[\tan \theta= 3 \cot \theta\] now \[\frac{ \tan \theta }{ \cot \theta }=3\] then \[\tan^2 \theta= 3-----and \tan \theta= \sqrt (3)\] find theta

OpenStudy (anonymous):

pi/3 correct

OpenStudy (ghazi):

correct

OpenStudy (zehanz):

\[\tan \theta= -\sqrt{3}\] is also possible...

OpenStudy (ghazi):

no , quantity under root cant be negative else it will be complex

OpenStudy (anonymous):

it is a plus or minus because of the square root. pi/3 +pi k, 2pi/3 +pi k

OpenStudy (anonymous):

Both of u were so much help thank u!

OpenStudy (ghazi):

yea you're right sorry i am bit sleepy